OpenStudy (anonymous):

Solve: sec²θ = 2tanθ+4 over the interval [0,2pi)

OpenStudy (yuki):

sec^2 = 1 + tan^2 so if you substitute that you will get \[\tan^2(x) -2\tan(x) -3 = 0\]

OpenStudy (yuki):

if you let tan(X) = a, it's obvious that it's a quadratic such that (a-3)(a-2) = 0 so,

OpenStudy (yuki):

\[\tan(x) = 3,-1\]

OpenStudy (anonymous):

would it be (a+1)(a-3) ?

OpenStudy (anonymous):

sorry, you did get that

OpenStudy (yuki):

there are infinitely many solutions for x, but since the interval is only from 0 to 2pi, x = \[\tan^{-1}(3), \tan^{-1}(3) + \pi , \tan^{-1}(-1)\]

OpenStudy (yuki):

you need a calculator for arctan(3) but arctan(-1) is a value we know which is -45 degrees. in terms of radians in the 3rd quadrant, it would be \[7\pi \over 4\]

OpenStudy (yuki):

got it ? :)

OpenStudy (anonymous):

so is the answer 7pi/4 and that's it?

OpenStudy (yuki):

no, arctan(3) and arctan(3)+pi are also answers

OpenStudy (yuki):

it's just a number that we cannot say exactly what it is

OpenStudy (anonymous):

oh, because my answer has to be in [0,2pi) form.....no decimals

OpenStudy (yuki):

Iris, the exact answer IS arctan(3), arctan(3) + pi, 7pi/4